# Enhancing Code-Switching Speech Recognition with LID-Based Collaborative Mixture of Experts Model

### v1

@@ 3.1 Dense-MoE && Sparse-MoE
Figure 1 illustrates two types of MoE models. Subfigure (a) depicts Dense-MoE, which was the original form of MoE when it was first proposed [27]. In Dense-MoE, all expert networks participate in forward propagation and are weighted and fused according to the weights obtained from gating networks to obtain more comprehensive features. The previously mentioned Bi-Encoder architecture [19] can be considered a variant of Dense-MoE because all expert networks are involved in the computation. Subfigure (b) represents Sparse-MoE models, where only a tiny portion of expert networks are typically selected to participate in computation. Sparse-MoE has recently gained popularity in large language models because it can increase the model parameter size while maintaining constant inference efficiency.[26].
@@ 3.2 LID-driven expert routing mechanism
$\boldsymbol{P}=\frac{\exp\left(\frac{h_{{\text{lid}}}^{j}}{T}\right)}{\sum_{j=1}^{3}\exp\left(\frac{h_{\text{lid}}^{j}}{T}\right)}$ (6)
@@ 4.3 Results of proposed model and baselines
System Model Params GFLOPs $\downarrow$ RTF $\downarrow$ CN EN CS Average
@@
Our proposed model $S8$ outperforms previous bi-encoder methods $S3$, $S4$, and $S5$ while significantly reducing GFLOPs and RTF. This improvement is mainly due to its compact topology, resembling a single encoder, and benefiting from the efficient inference of sparse MoE. Compared to the $S1$, we achieve an average relative performance improvement of 10.17% across the three test sets, with relative performance improvements of 22.7%, 21.6%, and 6% on CN, EN, and CS, respectively, while maintaining comparable computational costs.
@@ 4.4 Ablation study
We also experimented with removing the CS group, forcing the model to choose between the Chinese and English groups. The results show a significant performance degradation in CS scenarios and no advantage in CN and EN scenarios. In this case, the LID weights did not function effectively, and the lack of inter-group collaboration could cause this impact. These experiments also demonstrate that if we want the model to focus on a particular language, we can allocate more experts to the expert group of that language. This flexible design approach is challenging to achieve with other sparse MoEs that are not based on LID.

### v2

@@ 3.1 Dense-MoE && Sparse-MoE
Fig.1 illustrates two types of MoE models. Subfigure (a) depicts Dense-MoE, which was the original form of MoE when it was first proposed [27]. In Dense-MoE, all expert networks participate in forward propagation and are weighted and fused according to the weights obtained from gating networks to obtain more comprehensive features. The previously mentioned Bi-Encoder architecture [19] can be considered a variant of Dense-MoE because all expert networks are involved in the computation. Subfigure (b) represents Sparse-MoE models, where only a tiny portion of expert networks are typically selected to participate in computation. Sparse-MoE has recently gained popularity in large language models because it can increase the model parameter size while maintaining constant inference efficiency[26].
@@ 3.2 LID-driven expert routing mechanism
$\boldsymbol{P}=\frac{\exp\left(h_{\text{lid}}^{j}/T\right)}{\sum_{j=1}^{3}\exp\left(h_{\text{lid}}^{j}/T\right)}$ (6)
@@ 4.3 Results of proposed model and baselines
System Model Params FLOPs $\downarrow$ RTF $\downarrow$ CN EN CS Average
@@
As shown in Table 2, our proposed model $S8$ outperforms previous bi-encoder methods $S3$, $S4$, and $S5$ while significantly reducing GFLOPs and RTF. This improvement is mainly due to its compact topology, resembling a single encoder, and benefiting from the efficient inference of sparse MoE. Compared to the $S1$, we achieve an average relative performance improvement of 10.17% across the three test sets, with relative performance improvements of 22.7%, 21.6%, and 6% on CN, EN, and CS, respectively, while maintaining comparable computational costs.
@@ 4.4 Ablation study
We also experimented with removing the CS group, forcing the model to choose between the Chinese and English groups. The results show a significant performance degradation in CS scenarios and no advantage in CN and EN scenarios. In this case, the LID weights did not function effectively, and the lack of inter-group collaboration could cause this impact. These experiments also demonstrate that if we want the model to focus on a particular language, we can allocate more experts to the expert group of that language. This flexible design approach is challenging for other sparse MoEs that are not based on LID.